OpenStudy (kiki99):

What is the exact value of x in the exponential equation 25.3 + e7x = 97.3?

OpenStudy (kiki99):

@jabez177

OpenStudy (kiki99):

e^7x*

OpenStudy (alekos):

start by subtracting 25.3 from both sides

OpenStudy (alekos):

what do you get?

OpenStudy (kiki99):

e^7x+72

OpenStudy (alekos):

No, e^7x = 72 correct?

OpenStudy (kiki99):

oh ok

OpenStudy (kiki99):

whats next

OpenStudy (alekos):

Now take the natural log of both sides

OpenStudy (kiki99):

i forgot how to do that

OpenStudy (alekos):

Just take the Natural Log of both sides...... Lne^7x = Ln72 OK?

OpenStudy (kiki99):

how do i put the e in the calculator

OpenStudy (kiki99):

it give me error

OpenStudy (alekos):

Not yet, you're jumping the gun. You have to use one of the log rules on the LHS Ln(a^b) = bLna

OpenStudy (kiki99):

im lost i dont remember learning that last part. i dont know what im doing

OpenStudy (alekos):

How about this Log(a^b) = b Loga

OpenStudy (kiki99):

x = ln 72 − 7 x=In10.3/7 x=In72/7 x=In72/In7 are the choices

OpenStudy (alekos):

Yes, its definitely one of those choices.

OpenStudy (alekos):

LHS = Ln(e^7x) = ?

OpenStudy (alekos):

use what i've given you

OpenStudy (kiki99):

okay hold on

OpenStudy (kiki99):

I Dont Know What to do.

OpenStudy (kiki99):

i need to find the value of x not solve the equation

OpenStudy (alekos):

If Ln(a^b) = bLna then what does Ln(e^7x) = ?

OpenStudy (alekos):

In order to find the value of x you first need to re-arrange the equation to make x the subject

OpenStudy (alekos):

Goodnight